### 09 Mar 2026 v1

# Weighted Chernoff information and optimal loss exponent in context-sensitive hypothesis testing

Abstract
We consider context-sensitive (binary) hypothesis testing for i.i.d. observations under a multiplicative weight function. We establish the logarithmic asymptotic, as the sample size grows, of the optimal total loss (sum of type-I and type-II losses) and express the corresponding error exponent through a weighted Chernoff information between the competing distributions. Our approach embeds weighted geometric mixtures into an exponential family and identifies the exponent as the maximizer of its log-normaliser. We also provide concentration bounds for a tilted weighted log-likelihood and derive explicit expressions for Gaussian and Poisson models, as well as further parametric examples.
Key words and phrases: hypothesis testing, weighted Chernoff information, exponential family, error exponent, context-sensitive loss.
2020 Mathematics Subject Classifications: 62F03, 60F10.
1 Introduction
Let $\cal X$} be a Polish space with its Borel $\sigma$-algebra and let $X_{1}^{n}=(X_{1},\ldots,X_{n})$ be i.i.d. $\cal X$}-valued observations. We consider the simple hypotheses
$H_{0}:\ X_{1}^{n}\sim\mathbb{P}^{\otimes n}\qquad\text{versus}\qquad H_{1}:\ X_{1}^{n}\sim\mathbb{Q}^{\otimes n},$
where $\mathbb{P}$ and $\mathbb{Q}$ are probability measures on $\cal X$} dominated by a reference measure $\mu$. Without loss of generality one may take $\mu=\frac{1}{2}(\mathbb{P}+\mathbb{Q})$ and write $p=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mu}$ and $q=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mu}$. In the unweighted setting, the optimal sum of type-I and type-II error probabilities is characterized by $\mathrm{TV}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n})$ and can be written as
$\int\limits_{{\mathcal{X}}^{n}}\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),\penalty 10000\ \penalty 10000\ p(x_{1}^{n})=\prod\limits_{i=1}^{n}p(x_{i}),\penalty 10000\ \penalty 10000\ q(x_{1}^{n})=\prod\limits_{i=1}^{n}q(x_{i}).$ (1.1)
In the standard (unweighted) Bayesian setting, the decay rate of the optimal total error probability is governed by the Chernoff information [3, 10]:
$\rho_{\alpha}(p,q)$ $:=\int\limits_{\cal X}p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x),\qquad\alpha\in[0,1],$
$\rho(p,q)$ $:=\inf_{\alpha\in[0,1]}\rho_{\alpha}(p,q),$
$D_{C}(\mathbb{P},\mathbb{Q})$ $:=-\ln\rho(p,q)=\max_{\alpha\in[0,1]}\bigl[-\ln\rho_{\alpha}(p,q)\bigr].$ (1.2)
Here $\rho_{\alpha}$ is usually called the $\alpha$-skewed Bhattacharyya affinity coefficient, and $\rho(p,q)=\inf\limits_{\alpha\in[0,1]}\rho_{\alpha}(p,q)$ is the affinity coefficient. In view of Hölder’s inequality, $\rho_{\alpha}(p,q)\in[0,1]$.
Chernoff also introduced an asymptotic efficiency notion for comparing two experimental designs $e=\frac{\ln\rho_{1}}{\ln\rho_{2}}$ such that $n$ observations on one test are equivalent (i.e., give asymptotically the same total loss as $n\to\infty$) to $en$ observations on another test; see [3].
The focus of this paper is on a context-sensitive (weighted) analogue of this quantity and its logarithmic asymptotics as $n\to\infty$, cf. [5, 6]. In the context-sensitive setting, a nonnegative weight function (WF) $\varphi(x_{1}^{n})$ is given that reweights the loss of a wrong decision depending on the realised sample. It indicates the importance (or irrelevance) of the sample $X_{1}^{n}$ for the given statistical problem.
Throughout the paper, we assume that the weight is compatible with the i.i.d. structure and factorises across observations (with a slight abuse of notation, we use the same symbol $\varphi$ for the one-step weight and its product extension).
Assumption 1.1 (Factorised weight).
The weight function $\varphi({x}_{1}^{n})$ satisfies
$\varphi({x}_{1}^{n})=\prod_{i=1}^{n}\varphi(x_{i}),\qquad\varphi\geq 0.$ (1.3)
Assumption 1.1 is crucial: it yields single-letter weighted affinities $\rho^{\mathrm{w}}_{\alpha}(p,q)=\int_{\cal X}\varphi(x)p(x)^{\alpha}q(x)^{1-\alpha}\mathrm{d}\mu(x)$ and leads to an additive error exponent in the logarithmic asymptotics.
Intuitively, in the Bayesian framework with priors $\pi_{0}=\pi_{1}=\frac{1}{2}$ [^1] Here $\pi_{0},\pi_{1}$ denote the prior probabilities of $H_{0},H_{1}$., the Bayesian risk equals one half of the total loss. In particular, for one observation the weighted (Bayesian) risk can be written as
$\frac{1}{2}\int\limits_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x).$ (1.4)
Since $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}$ for every $\alpha\in[0,1]$, we obtain for any $\alpha$
$\int\limits_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x)\leq\int\limits_{\cal X}\varphi(x)p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)=\rho^{\mathrm{w}}_{\alpha}(p,q),$
and hence $\int\varphi\min\{p,q\}\,\mathrm{d}\mu\leq\inf\limits_{\alpha\in[0,1]}\rho^{\mathrm{w}}_{\alpha}(p,q).$ Thus, one obtains the upper bound
$\int\limits_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x)\leq\exp\{-D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})\},$
where $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q}):=\max\limits_{\alpha\in[0,1]}\bigl[-\ln\rho^{\mathrm{w}}_{\alpha}(p,q)\bigr]$ (see Definition 2.4 below). Under Assumption 1.1, the preceding bound factorises over $n$ i.i.d. observations and yields an upper bound of exponential order $\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})\}$ (up to a constant factor). Theorem 3.1 shows that this exponential rate is sharp on the logarithmic scale.
Main result and contributions.
Denote by $L_{n}^{*}$ the optimal total context-sensitive loss (sum of type-I and type-II losses, optimised over decision rules) for $n$ i.i.d. observations under Assumption 1.1. Our main theorem (Theorem 3.1) establishes an exact single-letter logarithmic asymptotic:
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\},\qquad n\to\infty,$ (1.5)
where the exponent is the weighted Chernoff information
$D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}\left[-\ln\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)\right].$ (1.6)
In particular, when $\varphi\equiv 1$, (1.6) reduces to the classical Chernoff information.
We also extend the exponent characterisation to a finite family of simple hypotheses: the optimal $M$-ary exponent is governed by the minimum pairwise weighted Chernoff information (cf. [9] for the corresponding unweighted principle). A key ingredient of the proofs is an exponential-family representation of weighted geometric mixtures: we embed the path $\alpha\mapsto\varphi\,p^{\alpha}q^{1-\alpha}$ into a likelihood-ratio exponential family and identify the Chernoff exponent as a maximiser of its log-normaliser. This viewpoint is closely related to the unweighted treatment via likelihood-ratio exponential families and log-normalisers in [7], while the present work may be viewed as a weighted extension in which the base measure is tilted by the context weight. In addition, we derive non-asymptotic concentration bounds for tilted weighted log-likelihood ratios, and provide explicit closed-form expressions for $D_{C}^{\mathrm{w}}$ in several parametric models (see Section 4).
Related work.
Exponential error rates in hypothesis testing originate in Chernoff’s seminal work [3]; see also Hoeffding [10]. The context-sensitive (weighted) testing framework and the weighted information measures used here are developed in [5, 6]. For $\varphi\equiv 1$, Chernoff information admits an information-geometric interpretation via likelihood-ratio exponential families generated by geometric mixtures and their log-normalisers, see [7]. Our results may be viewed as a weighted extension in which the base measure is tilted by the context weight $\varphi$. For background on information geometry of exponential families, see [11]. Multiple hypothesis testing and the principle that the optimal exponent is governed by the minimum pairwise Chernoff information are discussed in [9].
Structure of the paper.
Section 2 gives formal definitions of the weighted Bhattacharyya affinities and the weighted Chernoff information used throughout (Subsection 2.2; see Definitions 2.32.4). Section 3 states and proves the main asymptotic result (1.5) and develops the exponential-family/log-normaliser representation together with its information-geometric consequences. Subsection 3.4 studies the tilted weighted log-likelihood and derives non-asymptotic concentration bounds. Section 4 illustrates the general theory on Gaussian, Poisson, and exponential models, and includes an application to $M$-ary hypothesis testing (Subsection 4.4). Some auxiliary results are collected in the appendices.
2 Problem set-up and weighted divergences
2.1 Context-sensitive losses and weighted total variation
We keep the binary i.i.d. model from Section 1 and work under Assumption 1.1. In particular,
$p(x_{1}^{n})=\prod_{i=1}^{n}p(x_{i}),\qquad q(x_{1}^{n})=\prod_{i=1}^{n}q(x_{i}),\qquad\varphi(x_{1}^{n})=\prod_{i=1}^{n}\varphi(x_{i}).$
Define the $\varphi$-tilted (reweighted) densities
$p^{*}(x):=\frac{\varphi(x)p(x)}{E_{\varphi}(p)},\qquad q^{*}(x):=\frac{\varphi(x)q(x)}{E_{\varphi}(q)},\qquad E_{\varphi}(p):=\int\varphi(x)p(x)\,d\mu(x),$
(and similarly for $E_{\varphi}(q)$). Throughout this section we assume that $E_{\varphi}(p),E_{\varphi}(q)\in(0,\infty)$. (Equivalently, $\rho^{\mathrm{w}}_{0}(p,q),\rho^{\mathrm{w}}_{1}(p,q)\in(0,\infty).$) Then $p^{*},q^{*}$ are probability densities and, under $\varphi(x_{1}^{n})=\prod\limits_{i=1}^{n}\varphi(x_{i})$, we have $\varphi(x_{1}^{n})p(x_{1}^{n})=E_{\varphi}(p)^{n}(p^{*})^{\otimes n}(x_{1}^{n})$ (and similarly for $q$).
Under Assumption 1.1, we have
$\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,p(x_{1}^{n})\,d\mu^{\otimes n}=(E_{\varphi}(p))^{n},\qquad\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,q(x_{1}^{n})\,d\mu^{\otimes n}=(E_{\varphi}(q))^{n}.$
Let $\mathcal{D}$ denote the class of (possibly randomised) decision rules $D:{\cal X}^{n}\to[0,1]$, where $D(x_{1}^{n})$ is the probability of deciding in favour of $H_{1}$ after observing $x_{1}^{n}$. (Deterministic rules correspond to $D\in\{0,1\}$.)
For $D\in\mathcal{D}$, define the context-sensitive type-I and type-II losses by
$\alpha_{\varphi}(D)$ $:=\mathbb{E}_{\mathbb{P}^{\otimes n}}\!\big[\varphi(X_{1}^{n})\,D(X_{1}^{n})\big]=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,D(x_{1}^{n})\,p(x_{1}^{n})\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (2.1)
$\beta_{\varphi}(D)$ $:=\mathbb{E}_{\mathbb{Q}^{\otimes n}}\!\big[\varphi(X_{1}^{n})\,(1-D(X_{1}^{n}))\big]=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,(1-D(x_{1}^{n}))\,q(x_{1}^{n})\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (2.2)
and the corresponding total loss
$L_{n}(D):=\alpha_{\varphi}(D)+\beta_{\varphi}(D),\qquad L_{n}^{*}:=\inf_{D\in\mathcal{D}}L_{n}(D).$
Proposition 2.1 (Pointwise form of the optimal total loss).
For each $n\geq 1$,
$L_{n}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$ (2.3)
Moreover, an optimal (deterministic) decision rule is given by the likelihood-ratio test
$D_{n}^{*}(x_{1}^{n})={\bf 1}\{q(x_{1}^{n})\geq p(x_{1}^{n})\}$
(with any measurable tie-breaking on $\{p=q\}$).
Proof.
Fix $x_{1}^{n}$. The integrand in $L_{n}(D)$ equals
$\varphi(x_{1}^{n})\big(p(x_{1}^{n})D(x_{1}^{n})+q(x_{1}^{n})(1-D(x_{1}^{n}))\big)=\varphi(x_{1}^{n})\big(q(x_{1}^{n})+D(x_{1}^{n})(p(x_{1}^{n})-q(x_{1}^{n}))\big).$
Minimising pointwise over $D(x_{1}^{n})\in[0,1]$ yields $D_{n}^{*}(x_{1}^{n})=1$ when $p(x_{1}^{n})\leq q(x_{1}^{n})$ and $D_{n}^{*}(x_{1}^{n})=0$ when $p(x_{1}^{n})>q(x_{1}^{n})$, giving (2.3). ∎
We also use the weighted total variation distance
${\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n}):=\frac{1}{2}\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,\big|p(x_{1}^{n})-q(x_{1}^{n})\big|\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$ (2.4)
Remark 2.2.
For $\varphi\equiv 1$ this reduces to the usual total variation distance. If $\varphi$ vanishes on a non-negligible set, ${\rm TV}_{\varphi}$ is, in general, a pseudo-distance; this is sufficient for our purposes since it characterises the weighted losses.
Using $\min\{a,b\}=\frac{1}{2}(a+b-|a-b|)$ in (2.3) and the definition of ${\rm TV}_{\varphi}$ yields
$L_{n}^{*}=\frac{1}{2}\Big((E_{\varphi}(p))^{n}+(E_{\varphi}(q))^{n}\Big)-{\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n}).$ (2.5)
2.2 Weighted affinities and Chernoff information
We introduce the weighted Bhattacharyya affinities and the weighted Chernoff information. Assume that $\rho^{\mathrm{w}}_{\alpha}(p,q)\in(0,\infty)$ for all $\alpha\in[0,1]$.
Definition 2.3 (Weighted Bhattacharyya coefficient and distance).
For $\alpha\in[0,1]$ define the weighted $\alpha$-skewed Bhattacharyya affinity coefficient
$\rho^{\mathrm{w}}_{\alpha}(p,q):=\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x),$ (2.6)
and the corresponding weighted Bhattacharyya distance
$D^{\mathrm{w}}_{B,\alpha}(p,q):=-\ln\rho^{\mathrm{w}}_{\alpha}(p,q).$ (2.7)
Definition 2.4 (Weighted Chernoff information).
The weighted Chernoff information divergence between $\mathbb{P}$ and $\mathbb{Q}$ is
$D^{\mathrm{w}}_{C}({\mathbb{P}},{\mathbb{Q}})=\max_{\alpha\in[0,1]}\left[-\ln\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)\right]=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(p,q).$ (2.8)
A maximiser $\alpha^{*}=\alpha^{*}(p,q)$ in (2.8) is called the optimal Chernoff parameter.
Remark 2.5.
The weighted Chernoff information is symmetric: $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=D_{C}^{\mathrm{w}}(\mathbb{Q},\mathbb{P})$, since $\rho^{\mathrm{w}}_{\alpha}(p,q)=\rho^{\mathrm{w}}_{1-\alpha}(q,p)$. In general, however, $D_{C}^{\mathrm{w}}$ does not satisfy the triangle inequality and is therefore a divergence rather than a metric.
Remark 2.6.
Under Assumption 1.1, for every $\alpha\in[0,1]$,
$\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,p(x_{1}^{n})^{\alpha}q(x_{1}^{n})^{1-\alpha}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n})=\bigl(\rho^{\mathrm{w}}_{\alpha}(p,q)\bigr)^{n}.$
Consequently, the weighted Bhattacharyya distances are additive in $n$ and the corresponding Chernoff exponent is of single-letter form.
3 Main results
3.1 Asymptotics of the optimal sum of losses
Recall from Section 2 that
$L_{n}^{*}:=\inf_{D\in\mathcal{D}}\bigl[\alpha_{\varphi}(D)+\beta_{\varphi}(D)\bigr],$
and that Proposition 2.1 yields (2.3). The next theorem identifies its exact logarithmic asymptotic rate with the weighted Chernoff information from Definition 2.4.
Theorem 3.1 (Optimal sum of context-sensitive losses).
Consider the binary hypotheses $H_{0}:\ X_{1}^{n}\sim\mathbb{P}^{\otimes n}$ versus $H_{1}:\ X_{1}^{n}\sim\mathbb{Q}^{\otimes n}$, under Assumption 1.1. Let $\mathcal{D}$ be the class of (possibly randomised) decision rules $D:{\cal X}^{n}\to[0,1]$, and let $\alpha_{\varphi}(D)$, $\beta_{\varphi}(D)$ be defined by (2.1)–(2.2). Assume that $p,q>0$ $\mu$-a.e. and that $\rho^{\mathrm{w}}_{\alpha}(p,q)\in(0,\infty)$ for all $\alpha\in[0,1]$.
Then, as $n\to\infty$,
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}.$ (3.1)
Equivalently,
$\lim_{n\to\infty}-\frac{1}{n}\ln L_{n}^{*}=D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q}).$
Proof.
By Proposition 2.1,
$L_{n}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$
For any $\alpha\in[0,1]$, $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}$, hence by factorisation,
$L_{n}^{*}\leq\bigl(\rho^{\mathrm{w}}_{\alpha}(p,q)\bigr)^{n}.$
Taking $\inf_{\alpha\in[0,1]}$ yields $\liminf_{n\to\infty}-\frac{1}{n}\ln L_{n}^{*}\geq D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$.
Fix $\alpha^{*}\in\arg\min_{\alpha\in[0,1]}\rho^{\mathrm{w}}_{\alpha}(p,q)$ and set $r_{\alpha^{*}}(x):=\varphi(x)p(x)^{\alpha^{*}}q(x)^{1-\alpha^{*}}/\rho^{\mathrm{w}}_{\alpha^{*}}(p,q)$. Let $S_{n}=\sum\limits_{i=1}^{n}\ln\frac{p(X_{i})}{q(X_{i})}$. A change of measure gives
$\begin{array}[]{l}L_{n}^{*}=(\rho^{*}_{\alpha^{*}}(p,q))^{n}\mathbb{E}_{{r^{*}}_{\alpha^{*}}^{\otimes n}}\Big[e^{-\alpha^{*}S_{n}}\mathbf{1}(S_{n}<0)+e^{(1-\alpha^{*})S_{n}}\mathbf{1}(S_{n}\geq 0)\Big].\end{array}$ (3.2)
The bracket is bounded by $1$.
Let $F(\alpha):=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$. Under the regularity assumptions of [5], $m:=\mathbb{E}_{r_{\alpha^{*}}}[\ln(p(X)/q(X))]=F^{\prime}(\alpha^{*})$ exists, and since $\alpha^{*}$ minimises $F$ on $[0,1]$, we have $m=0$ if $\alpha^{*}\in(0,1)$, while $m\leq 0$ if $\alpha^{*}=0$ and $m\geq 0$ if $\alpha^{*}=1$. Hence $S_{n}/n\to m$ in $r_{\alpha^{*}}$-probability. For any $\varepsilon>0$, on the event $\{|S_{n}/n-m|\leq\varepsilon\}$ the bracket is $\geq e^{-\varepsilon n}$, while the event has probability 1o(1)1-o(1). Therefore the expectation in brackets equals $\exp\{o(n)\}$.
Hence $L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}$. For a more detailed treatment under additional regularity conditions we refer to [5]. ∎
Corollary 3.2 (Asymptotics of the weighted total variation).
Under the assumptions of Theorem 3.1, the weighted total variation satisfies
${\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n})=\frac{1}{2}\Big((E_{\varphi}(p))^{n}+(E_{\varphi}(q))^{n}\Big)-\exp\{-nD^{\mathrm{w}}_{C}(\mathbb{P},\mathbb{Q})+o(n)\},\penalty 10000\ \penalty 10000\ n\to\infty,$ (3.3)
where $E_{\varphi}(r)=\int_{\cal X}\varphi(x)r(x)\,\mathrm{d}\mu(x)$.
Proof.
Combine the identity (2.5) with (3.1). ∎
Remark $3.$
The weighted $\alpha$-skewed Bhattacharyya distance (2.7) appears in many papers, see e.g. [7]. Definition 2.4 shows that the weighted Chernoff information divergence is the maximally skewed weighted Bhattacharyya distance.
3.2 Exponential-family representation and uniqueness of $\alpha^{*}$
In order to develop an effective computational procedure and to connect the weighted Chernoff information to information geometry, we embed the weighted geometric mixtures of $p$ and $q$ into a one-parameter likelihood-ratio exponential family. For $\alpha\in[0,1]$ define
$Z_{pq}(\alpha):=\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)=\rho^{\mathrm{w}}_{\alpha}(p,q),$ (3.4)
and the corresponding normalised density
${\cal E}_{pq}=\left\{(pq)_{\alpha}(x):=\frac{\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}}{Z_{pq}(\alpha)}\;:\;\alpha\in[0,1]\right\}.$ (3.5)
By assumption $Z_{pq}(\alpha)\in(0,\infty)$, so $(pq)_{\alpha}$ is well-defined as a probability density w.r.t. $\mu$.
Set $t(x):=\ln\frac{p(x)}{q(x)}$ and $k_{pq}(x):=\ln\varphi(x)+\ln q(x)$. Then $(pq)_{\alpha}$ admits the exponential-family form
$(pq)_{\alpha}(x)$ $=\exp\!\left\{\alpha\,t(x)-F_{pq}(\alpha)+k_{pq}(x)\right\},$ (3.6)
$F_{pq}(\alpha):$ $=\ln Z_{pq}(\alpha)=-D^{\mathrm{w}}_{B,\alpha}(p,q).$ (3.7)
In particular, $t(X)$ is a sufficient statistic for the family ${\cal E}_{pq}$.
The log-normaliser $F_{pq}$ is convex on $[0,1]$; if $\ln\frac{p}{q}$ is not $\mu$-a.e. constant on $\{\varphi>0\}$, then $F_{pq}$ is strictly convex and the maximiser $\alpha^{*}$ in (2.8) is unique. By Hölder’s inequality,
$Z_{pq}(\alpha)=\int_{\cal X}(\varphi p)^{\alpha}(\varphi q)^{1-\alpha}\,\mathrm{d}\mu\leq(E_{\varphi}(p))^{\alpha}(E_{\varphi}(q))^{1-\alpha}.$
Finally, note that $Z_{pq}(1)=E_{\varphi}(p)$ and $Z_{pq}(0)=E_{\varphi}(q)$, hence
$(pq)_{1}(x)=\frac{\varphi(x)p(x)}{E_{\varphi}(p)},\qquad(pq)_{0}(x)=\frac{\varphi(x)q(x)}{E_{\varphi}(q)},$
so ${\cal E}_{pq}$ is an exponential arc between the tilted versions of $\mathbb{P}$ and $\mathbb{Q}$. By the definition, the following identities hold:
$\begin{array}[]{l}D^{\mathrm{w}}_{C}(p,q)=D^{\mathrm{w}}_{B,\alpha^{*}(p,q)}(p,q)=D^{\mathrm{w}}_{B,\alpha^{*}(q,p)}(q,p)=D^{\mathrm{w}}_{C}(q,p).\end{array}$ (3.8)
3.3 Weighted Bregman divergence and information-geometric identities
Weighted KL divergence and weighted Bregman divergence.
This subsection collects information-geometric identities useful for analysing $\rho^{\mathrm{w}}_{\alpha}$ and for computing the optimal Chernoff parameter $\alpha^{*}$. We follow [5] for weighted Bregman divergences.
Let ${\cal E}=\{p_{\theta}:\theta\in\Theta\subset\mathbb{R}^{d}\}$ be a regular exponential family of densities (with respect to $\mu$),
$p_{\theta}(x)=\exp\{\theta^{\mathrm{T}}t(x)-F(\theta)+k(x)\},\qquad x\in{\cal X}.$ (3.9)
For a density $r$ set $E_{\varphi}(r):=\int_{\cal X}\varphi(x)\,r(x)\,\mathrm{d}\mu(x)$ and write $E_{\varphi}(\theta):=E_{\varphi}(p_{\theta})$.
Assume $E_{\varphi}(\theta)\in(0,\infty)$ for $\theta\in\Theta$ and define the tilted log-normaliser
$\hat{F}(\theta):=\ln\int_{\cal X}\varphi(x)\,e^{\theta^{\mathrm{T}}t(x)+k(x)}\,\mathrm{d}\mu(x)=F(\theta)+\ln E_{\varphi}(\theta).$ (3.10)
Equivalently, the tilted density is
$p_{\theta}^{*}(x)=\frac{\varphi(x)\,p_{\theta}(x)}{E_{\varphi}(\theta)}.$ (3.11)
Definition 3.4 (Weighted Kullback–Leibler divergence).
For densities $p,q$ on $\cal X$} define
$D^{\mathrm{w}}_{\rm KL}(p\|q):=\int_{\cal X}\varphi(x)\,p(x)\,\ln\frac{p(x)}{q(x)}\,\mathrm{d}\mu(x),$ (3.12)
whenever the integral is well defined in $(-\infty,\infty]$.
Definition 3.5 (Weighted Bregman divergence).
The weighted Bregman divergence associated with (F,F^)($\hat{F}$) is
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{2}):$ $=e^{\hat{F}(\theta_{2})-F(\theta_{2})}\Big[F(\theta_{1})-F(\theta_{2})-(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{2})\Big]$ (3.13)
$=E_{\varphi}(\theta_{2})\Big[F(\theta_{1})-F(\theta_{2})-(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{2})\Big].$
Remark 3.6.
For exponential families, the weighted KL divergence admits the Bregman representation $D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{1})$; see [5, Proposition 4.1].
Proposition 3.7 (Weighted KL as weighted Bregman divergence).
Assume that the integral in (3.12) is well-defined in $(-\infty,\infty]$. Then for any $\theta_{1},\theta_{2}\in\Theta$,
$D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{1}).$ (3.14)
Proof.
This identity is stated in [5, Proposition 4.1]; we give a short derivation for completeness. By (3.9),
$\ln\frac{p_{\theta_{1}}(x)}{p_{\theta_{2}}(x)}=(\theta_{1}-\theta_{2})^{\mathrm{T}}t(x)-(F(\theta_{1})-F(\theta_{2})).$
Substituting this into (3.12) yields
$D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=(\theta_{1}-\theta_{2})^{\mathrm{T}}\!\int_{\cal X}\varphi(x)t(x)p_{\theta_{1}}(x)\,\mathrm{d}\mu(x)-(F(\theta_{1})-F(\theta_{2}))\,E_{\varphi}(\theta_{1}).$
[…]
$\nabla\hat{F}(\theta_{1})=\frac{\int_{\cal X}\varphi(x)t(x)e^{\theta_{1}^{\mathrm{T}}t(x)+k(x)}\,\mathrm{d}\mu(x)}{\int_{\cal X}\varphi(x)e^{\theta_{1}^{\mathrm{T}}t(x)+k(x)}\,\mathrm{d}\mu(x)}=\frac{\int_{\cal X}\varphi(x)t(x)p_{\theta_{1}}(x)\,\mathrm{d}\mu(x)}{E_{\varphi}(\theta_{1})}.$
Hence $\int\varphi tp_{\theta_{1}}\,\mathrm{d}\mu=E_{\varphi}(\theta_{1})\nabla\hat{F}(\theta_{1})$, and therefore
$D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=E_{\varphi}(\theta_{1})\Big[F(\theta_{2})-F(\theta_{1})-(\theta_{2}-\theta_{1})^{\mathrm{T}}\nabla\hat{F}(\theta_{1})\Big]=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{1}),$
which proves (3.14). ∎
Proposition 3.8 (Primal–dual identities for (weighted) Bregman divergences).
Let $F$ be a log-normaliser of a regular exponential family and let $F^{*}$ denote its Legendre transform. Write $\theta^{*}=\nabla F(\theta)$ and $\theta=\nabla F^{*}(\theta^{*})$.
(a) Weighted one-parameter identity. Assume $d=1$ (one-parameter case) and let $\theta_{i}^{*}:=F^{\prime}(\theta_{i})$. Then the following weighted analogue of the classical Bregman duality holds:
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{2})=B^{\mathrm{w}}_{\varphi,F^{*}}(\theta_{2}^{*},\theta_{1}^{*})-(\theta_{1}-\theta_{2})\,\ln E_{\varphi}(\theta_{2})+(\theta_{2}^{*}-\theta_{1}^{*})\,\ln E_{\varphi}(\theta_{1}),$ (3.15)
where $B^{\mathrm{w}}_{\varphi,F}$ is as in Definition 3.5 and $B^{\mathrm{w}}_{\varphi,F^{*}}$ is defined analogously (with $F$ replaced by $F^{*}$ and with the convention $E_{\varphi}(\theta^{*}):=E_{\varphi}(\theta)$ under $\theta=\nabla F^{*}(\theta^{*})$).
(b) Classical identity. For any $d\geq 1$, the (unweighted) Bregman divergence admits the standard Legendre representation
$B_{F}(\theta_{0},\theta_{1})=F(\theta_{0})+F^{*}(\theta_{1}^{*})-\theta_{0}^{\mathrm{T}}\theta_{1}^{*},\qquad\theta_{1}^{*}=\nabla F(\theta_{1}),$ (3.16)
where $B_{F}(\theta_{0},\theta_{1})=F(\theta_{0})-F(\theta_{1})-(\theta_{0}-\theta_{1})^{\mathrm{T}}\nabla F(\theta_{1})$ is the usual (unweighted) Bregman divergence.
Proof.
Part (a) is a weighted extension of the classical duality $B_{F}(\theta_{1},\theta_{2})=B_{F^{*}}(\theta_{2}^{*},\theta_{1}^{*})$ and follows by combining the weighted representation (3.13) with Legendre relations; see also [5]. Part (b) is standard. ∎
Weighted Chernoff/Bhattacharyya quantities inside an exponential family.
Let $p_{\theta_{1}},p_{\theta_{2}}\in{\cal E}$ and $\theta_{\alpha}:=\alpha\theta_{1}+(1-\alpha)\theta_{2}$. A direct calculation yields
$\rho^{\mathrm{w}}_{\alpha}(p_{\theta_{1}},p_{\theta_{2}})$ $=\int_{\cal X}\varphi(x)\,p_{\theta_{1}}(x)^{\alpha}p_{\theta_{2}}(x)^{1-\alpha}\,\mathrm{d}\mu(x)$ (3.17)
$=\exp\{\hat{F}(\theta_{\alpha})-\alpha F(\theta_{1})-(1-\alpha)F(\theta_{2})\}.$
Consequently,
$D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})$ $=\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2})-\hat{F}(\theta_{\alpha})$ (3.18)
$=U_{F,\alpha}(\theta_{1},\theta_{2})-\ln E_{\varphi}(\theta_{\alpha}),$
where $U_{F,\alpha}(\theta_{1},\theta_{2}):=\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2})-F(\theta_{\alpha})$ is the (unweighted) Jensen/Burbea–Rao divergence induced by $F$. In particular, when $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and $D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})=U_{F,\alpha}(\theta_{1},\theta_{2})$.
Remark 3.9 (Geometric mixtures and tilting by $\varphi$).
In particular, when $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and the normalised geometric mixture $p_{\theta_{1}}^{\alpha}p_{\theta_{2}}^{1-\alpha}$ belongs to the same exponential family, namely $\propto p_{\theta_{\alpha}}$ with $\theta_{\alpha}=\alpha\theta_{1}+(1-\alpha)\theta_{2}$.
Proposition 3.10 (Optimal Chernoff parameter in an exponential family).
Assume that $\hat{F}$ is strictly convex on the segment $\theta_{1},\theta_{2}]$ and that the maximiser $\alpha^{*}\in(0,1)$ exists. Then $\alpha^{*}$ is unique and satisfies
$(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{\alpha^{*}})=F(\theta_{1})-F(\theta_{2}),\qquad\theta_{\alpha^{*}}=\alpha^{*}\theta_{1}+(1-\alpha^{*})\theta_{2},$ (3.19)
with $\nabla\hat{F}(\theta)=\mathbb{E}_{p_{\theta}^{*}}[t(X)]$.
Proof.
Differentiate (3.18) with respect to $\alpha$ and use $\frac{\mathrm{d}}{\mathrm{d}\alpha}\theta_{\alpha}=\theta_{1}-\theta_{2}$. Strict convexity of $\hat{F}$ on $\theta_{1},\theta_{2}]$ implies strict concavity of $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(p_{\theta_{1}},p_{\theta_{2}})$, hence uniqueness. ∎
Proposition 3.11 (Chernoff information as a Jensen-type divergence and a Bregman bisector).
Let $p_{\theta_{1}},p_{\theta_{2}}\in{\cal E}$ and assume that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 exists and is unique. Set $\theta_{\alpha}=\alpha\theta_{1}+(1-\alpha)\theta_{2}$. Then
$D_{C}^{\mathrm{w}}(p_{\theta_{1}},p_{\theta_{2}})=D^{\mathrm{w}}_{B,\alpha^{*}}(p_{\theta_{1}},p_{\theta_{2}})=\alpha^{*}F(\theta_{1})+(1-\alpha^{*})F(\theta_{2})-\hat{F}(\theta_{\alpha^{*}}).$ (3.20)
Moreover, $\theta_{\alpha^{*}}$ is characterised by the weighted Bregman bisector condition
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{\alpha^{*}}),$ (3.21)
and the common value recovers the Chernoff information as
$D_{C}^{\mathrm{w}}(p_{\theta_{1}},p_{\theta_{2}})$ $=\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})-\ln E_{\varphi}(\theta_{\alpha^{*}})$ (3.22)
$=\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{\alpha^{*}})-\ln E_{\varphi}(\theta_{\alpha^{*}}).$ (3.23)
In the special case $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and $E_{\varphi}(\theta)\equiv 1$, so that (3.20) reduces to the classical Jensen divergence induced by $F$ and (3.22) becomes $D_{C}(p_{\theta_{1}},p_{\theta_{2}})=B_{F}(\theta_{1},\theta_{\alpha^{*}})=B_{F}(\theta_{2},\theta_{\alpha^{*}})$.
Proof.
Equation (3.20) is (3.18) at $\alpha=\alpha^{*}$. For (3.21), expand $B^{\mathrm{w}}_{\varphi,F}(\theta_{i},\theta_{\alpha^{*}})=E_{\varphi}(\theta_{\alpha^{*}})\big[F(\theta_{i})-F(\theta_{\alpha^{*}})-(\theta_{i}-\theta_{\alpha^{*}})^{\mathrm{T}}\nabla\hat{F}(\theta_{\alpha^{*}})\big]$ and use (3.19) to see that the difference vanishes. Finally, substituting $\theta_{1}-\theta_{\alpha^{*}}=(1-\alpha^{*})(\theta_{1}-\theta_{2})$ into $B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})/E_{\varphi}(\theta_{\alpha^{*}})$ and using (3.19) gives
$\frac{1}{E_{\varphi}(\theta_{\alpha^{*}})}B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{\alpha^{*}})=\alpha^{*}F(\theta_{1})+(1-\alpha^{*})F(\theta_{2})-F(\theta_{\alpha^{*}}).$
Since $\hat{F}(\theta_{\alpha^{*}})=F(\theta_{\alpha^{*}})+\ln E_{\varphi}(\theta_{\alpha^{*}})$, this yields (3.22). ∎
Derivative and weighted KL.
Recall $F_{pq}(\alpha)=\ln Z_{pq}(\alpha)=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$. Assuming differentiation under the integral sign is justified, we have
$F_{pq}^{\prime}(\alpha)=\mathbb{E}_{(pq)_{\alpha}}\!\left[\ln\frac{p(X)}{q(X)}\right],$ (3.24)
where $(pq)_{\alpha}$ is the Chernoff-tilted density from (3.5). In particular,
$F_{pq}^{\prime}(1)=\frac{1}{E_{\varphi}(p)}\int_{\cal X}\varphi(x)\,p(x)\,\ln\frac{p(x)}{q(x)}\,\mathrm{d}\mu(x)=\frac{1}{E_{\varphi}(p)}D^{\mathrm{w}}_{\rm KL}(p\|q).$ (3.25)
Analogously, $F^{\prime}_{pq}(0)=-\frac{1}{E_{\varphi}(q)}D^{\mathrm{w}}_{\rm KL}(q\|p)$.
Chernoff–KL.
Lemma 3.12.
Let $\alpha^{*}$ be a maximiser in Definition 2.4 and assume that $\alpha^{*}\in(0,1)$ (so that $F_{pq}^{\prime}(\alpha^{*})=0$ below). Set $F_{pq}(\alpha):=\ln\rho^{\mathrm{w}}_{\alpha}(p,q),r_{\alpha}:=(pq)_{\alpha}.$ Then
$D^{\mathrm{w}}_{C}(p,q)=D_{\rm KL}\!\big(r_{\alpha^{*}}\,\|\,r_{1}\big)-\ln E_{\varphi}(p)=D_{\rm KL}\!\big(r_{\alpha^{*}}\,\|\,r_{0}\big)-\ln E_{\varphi}(q),$ (3.26)
where $r_{1}(x)=\varphi(x)p(x)/E_{\varphi}(p)$ and $r_{0}(x)=\varphi(x)q(x)/E_{\varphi}(q)$.
Here $D_{\rm KL}$ denotes the standard (unweighted) Kullback–Leibler divergence.
Proof.
A direct computation yields, for $\alpha\in[0,1]$,
$D_{\rm KL}(r_{\alpha}\|r_{1})$ $=-(1-\alpha)F^{\prime}_{pq}(\alpha)-F_{pq}(\alpha)+F_{pq}(1),$
$D_{\rm KL}(r_{\alpha}\|r_{0})$ $=\alpha F^{\prime}_{pq}(\alpha)-F_{pq}(\alpha)+F_{pq}(0).$
At $\alpha=\alpha^{*}$ we have $F^{\prime}_{pq}(\alpha^{*})=0$. Since $F_{pq}(1)=\ln E_{\varphi}(p)$ and $F_{pq}(0)=\ln E_{\varphi}(q)$, the claim follows from $D_{C}^{\mathrm{w}}(p,q)=-F_{pq}(\alpha^{*})$. ∎
Corollary 3.13 (Chernoff information as a Bregman divergence on the Chernoff arc).
Let $F_{pq}(\alpha):=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$ and define the one-dimensional Bregman divergence
$B_{F_{pq}}(a,b):=F_{pq}(a)-F_{pq}(b)-(a-b)F^{\prime}_{pq}(b).$
Assume that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 is interior, [It is always inferior because of convexity?] so that $F^{\prime}_{pq}(\alpha^{*})=0$. Then
$D_{C}^{\mathrm{w}}(p,q)=B_{F_{pq}}(1,\alpha^{*})-\ln E_{\varphi}(p)=B_{F_{pq}}(0,\alpha^{*})-\ln E_{\varphi}(q).$ (3.27)
Equivalently,
$B_{F_{pq}}(1,\alpha^{*})$ $=D_{\rm KL}(r_{\alpha^{*}}\|r_{1})=D_{C}^{\mathrm{w}}(p,q)+\ln E_{\varphi}(p),$
$B_{F_{pq}}(0,\alpha^{*})$ $=D_{\rm KL}(r_{\alpha^{*}}\|r_{0})=D_{C}^{\mathrm{w}}(p,q)+\ln E_{\varphi}(q),$
with $r_{\alpha}=(pq)_{\alpha}$, $r_{1}=\varphi p/E_{\varphi}(p)$ and $r_{0}=\varphi q/E_{\varphi}(q)$.
Proof.
In the Chernoff exponential family $\{r_{\alpha}\}$ with log-normalizer $F_{pq}$, the KL–Bregman identity gives $D_{\rm KL}(r_{\alpha^{*}}\|r_{1})=B_{F_{pq}}(1,\alpha^{*})$ and $D_{\rm KL}(r_{\alpha^{*}}\|r_{0})=B_{F_{pq}}(0,\alpha^{*})$. Since $F^{\prime}_{pq}(\alpha^{*})=0$ and $F_{pq}(1)=\ln E_{\varphi}(p)$, $F_{pq}(0)=\ln E_{\varphi}(q)$, we obtain
$B_{F_{pq}}(1,\alpha^{*})=-F_{pq}(\alpha^{*})+F_{pq}(1)=D_{C}^{\mathrm{w}}(p,q)+\ln E_{\varphi}(p),$
and similarly for 0. Substituting this into Lemma 3.12 yields (3.27). ∎
Remark 3.14 (One-parameter case).
Assume $d=1$, $\theta_{1}\neq\theta_{2}$, and that the maximiser $\alpha^{*}\in(0,1)$ in Definition 2.4 is interior. Assume moreover that $\hat{F}^{\prime}$ is strictly increasing on $\Theta$ and set ${\hat{G}}=(\hat{F}^{\prime})^{-1}$. Then (3.19) yields
$\alpha^{*}=\frac{1}{\theta_{1}-\theta_{2}}\left({\hat{G}}\left(\frac{F(\theta_{1})-F(\theta_{2})}{\theta_{1}-\theta_{2}}\right)-\theta_{2}\right).$ (3.28)
When $\varphi\equiv 1$ we have $\hat{F}\equiv F$ and (3.28) reduces to the classical formula.
To illustrate the general identities above, we provide explicit expressions for $D^{\mathrm{w}}_{B,\alpha}$ and $D^{\mathrm{w}}_{C}$ in several parametric settings (Gaussian, Poisson, and exponential), see Section 4.
3.4 Tilted weighted likelihood and concentration bounds
Although the optimal decision rule in the context-sensitive setting is still based on the usual likelihood ratio $q/p$ (cf. Section 3), it is convenient to consider the tilted likelihood ratio $q^{*}/p^{*}$ as well. The reason is that the context weight $\varphi$ cancels pointwise in $q^{*}/p^{*}$ and affects inference only through the normalisation constants $E_{\varphi}(p)$ and $E_{\varphi}(q)$. We record below (i) a large-deviation representation for $L^{*}/n$ via the cumulant generating functions $\psi_{\mathbb{P}},\psi_{\mathbb{Q}}$ and their Legendre transforms, and (ii) a finite-$n$ concentration bound for $L^{*}$ based on standard martingale inequalities (Azuma–Hoeffding).
For the tilted distributions, the log-likelihood takes the form
$\begin{array}[]{l}L^{*}({X}_{1}^{n})=L^{*}(X_{1},\ldots,X_{n})=\sum\limits_{i=1}^{n}\ln\frac{q^{*}(X_{i})}{p^{*}(X_{i})}\\ \hskip 38.41139pt=\sum\limits_{i=1}^{n}\ln\frac{q(X_{i})}{p(X_{i})}-n\ln E_{\varphi}(q)+n\ln E_{\varphi}(p).\end{array}$ (3.29)
Here $E_{\varphi}(p)=\int\varphi(x)\,p(x)\,\mathrm{d}\mu(x)$.
In particular, since
$\ln\frac{q^{*}(x)}{p^{*}(x)}=\ln\frac{q(x)}{p(x)}+\ln E_{\varphi}(p)-\ln E_{\varphi}(q),$
we may equivalently rewrite likelihood-ratio threshold rules in terms of $L^{*}$. For example,
$\sum_{i=1}^{n}\ln\frac{q(X_{i})}{p(X_{i})}\geq 0\quad\Longleftrightarrow\quad L^{*}(X_{1}^{n})\geq n\big(\ln E_{\varphi}(p)-\ln E_{\varphi}(q)\big).$
Thus $L^{*}$ is the usual log-likelihood ratio, shifted by a constant determined by the context weight $\varphi$.
The log of the moment generating function and its Legendre transform take the form
$\psi_{\mathbb{P}}(\alpha)$ $=\ln\mathbb{E}_{\mathbb{P}}\left[e^{\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]$
$=\ln\int_{\cal X}q(x)^{\alpha}p(x)^{1-\alpha}\,\mathrm{d}\mu(x)-\alpha\ln E_{\varphi}(q)+\alpha\ln E_{\varphi}(p),$ (3.30)
$I_{\mathbb{P}}(r)$ $=\sup\limits_{\alpha}\left[\alpha r-\psi_{\mathbb{P}}(\alpha)\right],$
where $\alpha$ ranges over the set $\{\alpha\in\mathbb{R}:\psi_{\mathbb{P}}(\alpha)<\infty\}$ (and similarly for $\psi_{\mathbb{Q}}$).
Similarly,
$\psi_{\mathbb{Q}}(\alpha):$ $=\ln\mathbb{E}_{\mathbb{Q}}\left[e^{\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]=\ln\mathbb{E}_{\mathbb{P}}\left[e^{\ln\frac{q(X)}{p(X)}+\alpha\ln\frac{q^{*}(X)}{p^{*}(X)}}\right]$ (3.31)
$=\ln\int_{\cal X}q(x)^{\alpha+1}p(x)^{-\alpha}\,\mathrm{d}\mu(x)-\alpha\ln E_{\varphi}(q)+\alpha\ln E_{\varphi}(p).$
This implies the relation of Legendre transforms
$I_{\mathbb{Q}}(r)=I_{\mathbb{P}}(r)-r+\ln E_{\varphi}(p)-\ln E_{\varphi}(q).$ (3.32)
In particular, $I_{\mathbb{P}}(0)$ may be treated as a natural weighted version of the Chernoff divergence between $q$ and $p$:
$I_{\mathbb{P}}(0)$ $=\sup\limits_{\alpha}\left[-\ln\int_{\cal X}q(x)^{\alpha}p(x)^{1-\alpha}\,\mathrm{d}\mu(x)+\alpha\ln E_{\varphi}(q)-\alpha\ln E_{\varphi}(p)\right]$
$=:{\hat{D}}^{\rm w}_{C}(q,p).$ (3.33)
Interpretation. The value $I_{\mathbb{P}}(0)$ is the Chernoff–Cramér exponent controlling the tail event $\{L^{*}/n\geq 0\}$ under $\mathbb{P}$, i.e. (under standard regularity assumptions) $\mathbb{P}(L^{*}(X_{1}^{n})\geq 0)\asymp e^{-nI_{\mathbb{P}}(0)}$. In the unweighted case $\varphi\equiv 1$ we have $E_{\varphi}(p)=E_{\varphi}(q)=1$ and $I_{\mathbb{P}}(0)$ reduces to the classical Chernoff information between $p$ and $q$. We also stress that ${\hat{D}}_{C}^{\rm w}(q,p)=I_{\mathbb{P}}(0)$ is a tilted-likelihood exponent and is distinct from the weighted Chernoff information $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$ from Definition 2.4, which governs the optimal sum of context-sensitive losses.
Non-asymptotic concentration via a Doob martingale.
The rate functions $I_{\mathbb{P}}$ and $I_{\mathbb{Q}}$ capture the exponential scale of deviations of $L^{*}/n$ as $n\to\infty$. To obtain explicit finite-$n$ bounds, we now apply a standard martingale method to $L^{*}$ under $\mathbb{Q}$.
Consider the filtration ${\cal F}_{k}=\sigma(X_{1},\ldots,X_{k})$ and define the random variables $\{U_{k},k=0,\ldots,n\}$ by
$U_{k}$ $=\mathbb{E}_{\mathbb{Q}}\left[L^{*}(X_{1},\ldots,X_{n})|{\cal F}_{k}\right]$
$=\sum\limits_{j=1}^{k}\ln\frac{q(X_{j})}{p(X_{j})}+(n-k)D_{\rm KL}(\mathbb{Q}\|\mathbb{P})-n\ln E_{\varphi}(q)+n\ln E_{\varphi}(p),$ (3.34)
where $D_{\rm KL}(\mathbb{Q}\|\mathbb{P})=\int q(x)\ln\frac{q(x)}{p(x)}\,\mathrm{d}\mu(x)$ stands for the (unweighted) Kullback–Leibler divergence of $\mathbb{Q}$ and $\mathbb{P}$. Then
$U_{k}-U_{k-1}=\ln\frac{q(X_{k})}{p(X_{k})}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P}),\qquad k=1,\ldots,n.$ (3.35)
Observe that $\{U_{k},{\cal F}_{k}\}$ is a martingale w.r.t. $\mathbb{Q}$.
On the next step we consider a particular case when $d<\infty$ and $\sigma^{2}<\infty$, where $d=\sup\limits_{x\in{\cal X}}\left|\ln\frac{q(x)}{p(x)}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P})\right|$ and
$\sigma^{2}$ $=\mathbb{E}_{\mathbb{Q}}\left[(U_{k}-U_{k-1})^{2}|{\cal F}_{k-1}\right]$
$=\int_{\cal X}q(x)\left(\ln\frac{q(x)}{p(x)}-D_{\rm KL}(\mathbb{Q}\|\mathbb{P})\right)^{2}\,\mathrm{d}\mu(x).$ (3.36)
In view of a refined Azuma–Hoeffding/McDiarmid inequality [1, 2],
${\mathbb{P}}_{\mathbb{Q}}\bigg(L^{*}(X_{1},\ldots,X_{n})>(D_{\rm KL}(\mathbb{Q}\|\mathbb{P})+\beta)n-(\ln E_{\varphi}(q)-\ln E_{\varphi}(p))n\bigg)$
$\leq\exp\left\{-nD\left(\frac{\delta+\gamma^{*}}{1+\gamma^{*}}\bigg\|\frac{\gamma^{*}}{1+\gamma^{*}}\right)\right\},$ (3.37)
where $\delta=\frac{\sigma^{2}}{d^{2}}$, $\gamma^{*}=\frac{\beta^{*}}{d}$, $\beta^{*}=\beta-\ln E_{\varphi}(q)+\ln E_{\varphi}(p)$ and $D(p\|q)=p\ln\frac{p}{q}+(1-p)\ln\frac{1-p}{1-q}$ stands for the Kullback–Leibler divergence between the two Bernoulli distributions (p,1p)(p,1-p) and (q,1q)(q,1-q).
We use the following modified version of Azuma–Hoeffding inequality, see [2].
Lemma 3.15.
Let $\{U_{k},{\cal F}_{k}\}$ be a discrete-time real-valued martingale. Assume that, for some constants $d,\sigma>0$, the following two requirements are satisfied a.s. for every $k\in\{1,\ldots,n\}:$
$\begin{array}[]{l}|U_{k}-U_{k-1}|\leq d,\\ {\rm Var}[U_{k}-U_{k-1}|{\cal F}_{k-1}]\leq\sigma^{2}.\end{array}$ (3.38)
Then, for every $\beta\geq 0$,
${\mathbb{P}}\left(|U_{n}-U_{0}|\geq\beta n\right)\leq 2\exp\left\{-nD\left(\frac{\delta+\gamma}{1+\gamma}\bigg\|\frac{\gamma}{1+\gamma}\right)\right\},$ (3.39)
where $\delta=\frac{\sigma^{2}}{d^{2}}$ and $\gamma=\frac{\beta}{d}$. (Note that $\delta\in[0,1]$ automatically whenever $|U_{k}-U_{k-1}|\leq d$ a.s.)
In particular, the one-sided bound $\mathbb{P}(U_{n}-U_{0}\geq\beta n)\leq\exp\{-nD(\cdot)\}$ holds.
Theorem 3.16.
Set
$\beta^{*}=\beta-\ln E_{\varphi}(p)+\ln E_{\varphi}(q),\qquad\gamma^{*}=\frac{\beta^{*}}{d}.$
Under conditions $d<\infty$, $\sigma^{2}<\infty$, and $\beta^{*}\geq 0$,
${\mathbb{P}}_{\mathbb{Q}}\big(L^{*}(X_{1},\ldots,X_{n})\geq\beta n\big)\leq\exp\left\{-nD\left(\frac{\delta+\gamma^{*}}{1+\gamma^{*}}\bigg\|\frac{\gamma^{*}}{1+\gamma^{*}}\right)\right\}.$ (3.40)
Lemma 3.15 is quoted from [2]. Theorem 3.16 is its direct application to the tilted log-likelihood ratio (3.29); the only dependence on the context weight $\varphi$ is through $E_{\varphi}(p)$ and $E_{\varphi}(q)$.
4 Examples and applications
The identities of Section 3 reduce the computation of $D^{\mathrm{w}}_{B,\alpha}$ and $D^{\mathrm{w}}_{C}$ to evaluating the single-letter weighted affinity $\rho^{\mathrm{w}}_{\alpha}(p,q)=\int\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)$ (and then optimising over $\alpha\in[0,1]$ when computing $D^{\mathrm{w}}_{C}$). We work this out for several standard parametric families (Gaussian, Poisson, and exponential), highlighting how the context weight $\varphi$ modifies the classical formulas. When $\varphi\equiv 1$, the expressions reduce to the usual (unweighted) Bhattacharyya/Chernoff quantities available in the literature (e.g. [7, 9]). More involved computations outside the exponential-family setting (such as the Cauchy location–scale family) are deferred to the appendices.
4.1 Gaussian models
Throughout this subsection, the reference measure is the Lebesgue measure on $\mathbb{R}^{d}$. We compute the weighted Bhattacharyya coefficient
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}^{d}}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}x,\qquad\alpha\in[0,1],$
together with the weighted Bhattacharyya distance $D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q}):=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ and the weighted Chernoff information $D^{\mathrm{w}}_{C}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})$ (Definition 2.4). Note that, unlike the unweighted case, $\rho^{\mathrm{w}}_{\alpha}$ is not restricted to $[0,1]$ and $D^{\mathrm{w}}_{B,\alpha}$ (hence also $D_{C}^{\mathrm{w}}$) may take negative values.
Example 4.1 (Gaussian weighted Bhattacharyya coefficient with exponential weight).
Let $\mathbb{P}=\mathcal{N}(\mu_{1},\Sigma_{1})$ and $\mathbb{Q}=\mathcal{N}(\mu_{2},\Sigma_{2})$ on $\mathbb{R}^{d}$, where $\Sigma_{1}\succ 0$ and $\Sigma_{2}\succ 0$, and let $\varphi(x)=e^{\gamma^{\mathrm{T}}x}$ for some $\gamma\in\mathbb{R}^{d}$. Denote by $p,q$ the corresponding densities. For $\alpha\in[0,1]$ define
$\Sigma_{\alpha}$ $:=\left(\alpha\Sigma_{1}^{-1}+(1-\alpha)\Sigma_{2}^{-1}\right)^{-1},$
$\widetilde{\mu}_{\alpha}$ $:=\Sigma_{\alpha}\Big(\alpha\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\Sigma_{2}^{-1}\mu_{2}+\gamma\Big).$
Then
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}^{d}}e^{\gamma^{\mathrm{T}}x}\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}x$
$=\frac{|\Sigma_{\alpha}|^{1/2}}{|\Sigma_{1}|^{\alpha/2}|\Sigma_{2}|^{(1-\alpha)/2}}\exp\!\left\{-\frac{1}{2}\Big(\alpha\mu_{1}^{\mathrm{T}}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\mathrm{T}}\Sigma_{2}^{-1}\mu_{2}-\widetilde{\mu}_{\alpha}^{\mathrm{T}}\Sigma_{\alpha}^{-1}\widetilde{\mu}_{\alpha}\Big)\right\}.$ (4.1)
Consequently,
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=$
$=\frac{1}{2}\left(\alpha\mu_{1}^{\mathrm{T}}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\mathrm{T}}\Sigma_{2}^{-1}\mu_{2}-\widetilde{\mu}_{\alpha}^{\mathrm{T}}\Sigma_{\alpha}^{-1}\widetilde{\mu}_{\alpha}+\ln\frac{|\Sigma_{1}|^{\alpha}|\Sigma_{2}|^{1-\alpha}}{|\Sigma_{\alpha}|}\right).$ (4.2)
In particular, setting $\gamma=0$ (i.e. $\varphi\equiv 1$) reduces (4.1) to the classical (unweighted) Gaussian Bhattacharyya distance; see, e.g., [7].
Corollary 4.2 (Common covariance).
In Example 4.1, assume $\Sigma_{1}=\Sigma_{2}=\Sigma\succ 0$ and keep the exponential weight $\varphi(x)=e^{\gamma^{\mathrm{T}}x}$. Set $\delta:=\mu_{1}-\mu_{2}$, $\|v\|_{\Sigma^{-1}}^{2}:=v^{\mathrm{T}}\Sigma^{-1}v$, and $\mu_{\alpha}:=\alpha\mu_{1}+(1-\alpha)\mu_{2}$. Then, for any $\alpha\in[0,1]$,
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ $=\int_{\mathbb{R}^{d}}e^{\gamma^{\mathrm{T}}x}\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}x$
$=\exp\!\left\{-\frac{\alpha(1-\alpha)}{2}\|\delta\|_{\Sigma^{-1}}^{2}+\gamma^{\mathrm{T}}\mu_{\alpha}+\frac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma\right\},$ (4.3)
and therefore
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\frac{\alpha(1-\alpha)}{2}\|\delta\|_{\Sigma^{-1}}^{2}-\gamma^{\mathrm{T}}\mu_{\alpha}-\frac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma.$ (4.4)
If $\mu_{1}\neq\mu_{2}$ and the unconstrained maximiser
$\widetilde{\alpha}=\frac{1}{2}-\frac{\gamma^{\mathrm{T}}\delta}{\|\delta\|_{\Sigma^{-1}}^{2}}$
belongs to $(0,1)$, then $\alpha^{*}=\widetilde{\alpha}$; otherwise the maximum over $\alpha\in[0,1]$ is attained at the nearest boundary point $\alpha^{*}\in\{0,1\}$. In all cases,
$D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=D^{\mathrm{w}}_{B,\alpha^{*}}(\mathbb{P},\mathbb{Q}).$
In particular, for $\gamma=0$ (i.e. $\varphi\equiv 1$) we recover $\alpha^{*}=1/2$ and $D_{C}(\mathbb{P},\mathbb{Q})=\|\delta\|_{\Sigma^{-1}}^{2}/8$.
Proof.
The expression for $\rho^{\mathrm{w}}_{\alpha}$ follows by simplifying Example 4.1 under $\Sigma_{1}=\Sigma_{2}=\Sigma$, which makes the determinant prefactor equal to $1$ and yields a Gaussian MGF term $\exp(\gamma^{\mathrm{T}}\mu_{\alpha}+\tfrac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma)$. The maximiser follows by differentiating (4.4) in $\alpha$. ∎
Choosing an exponential weight $\varphi(x)=e^{\gamma^{\mathrm{T}}x}$ corresponds to exponential tilting: for a Gaussian $X\sim\mathcal{N}(\mu,\Sigma)$ this tilting keeps the covariance and shifts the mean to $\mu+\Sigma\gamma$ (with normalisation factor $\exp(\gamma^{\mathrm{T}}\mu+\tfrac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma)$), which is why the weighted affinities remain available in closed form. In particular, the optimal Chernoff parameter is no longer forced to be $\alpha^{*}=1/2$ and, as (4.4) shows, sufficiently strong tilting can push the maximiser to the boundary $\alpha^{*}\in\{0,1\}$.
4.2 Poisson models
Example 4.3 (Poisson model with exponential weight).
Let ${\cal X}=\mathbb{N}_{0}=\{0,1,2,\dots\}$ and let $\mu$ be the counting measure on $\cal X$}. Fix two hypotheses $\mathbb{P}=\mathrm{Poi}(\lambda_{1})$ and $\mathbb{Q}=\mathrm{Poi}(\lambda_{2})$ with $\lambda_{1},\lambda_{2}>0$, and write $p=p_{\lambda_{1}}$ and $q=q_{\lambda_{2}}$. We still consider the exponential weight $\varphi_{\gamma}(k)=e^{\gamma k},\penalty 10000\ \gamma\in\mathbb{R}.$ (For $\gamma=0$ we recover the unweighted case $\varphi\equiv 1$.)
For $\alpha\in[0,1]$ set
$\lambda_{\alpha}:=\lambda_{1}^{\alpha}\lambda_{2}^{\,1-\alpha}.$
(a) Weighted Bhattacharyya coefficient and Chernoff arc. A direct summation gives
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ $=\sum_{k=0}^{\infty}\varphi_{\gamma}(k)\,p(k)^{\alpha}q(k)^{1-\alpha}=\exp\!\big\{-\alpha\lambda_{1}-(1-\alpha)\lambda_{2}+e^{\gamma}\lambda_{\alpha}\big\}.$ (4.5)
Hence, by Definition 2.3,
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\alpha\lambda_{1}+(1-\alpha)\lambda_{2}-e^{\gamma}\lambda_{\alpha}.$ (4.6)
Moreover, the normalised weighted geometric mixture (Chernoff-tilted density) from (3.5) takes the form
$(pq)_{\alpha}(k)=\frac{\varphi_{\gamma}(k)\,p(k)^{\alpha}q(k)^{1-\alpha}}{\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})}=\exp\!\{-e^{\gamma}\lambda_{\alpha}\}\,\frac{(e^{\gamma}\lambda_{\alpha})^{k}}{k!}=\mathrm{Poi}(e^{\gamma}\lambda_{\alpha}).$
(b) Optimal Chernoff parameter. If $\lambda_{1}\neq\lambda_{2}$, then $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})$ is strictly concave on $[0,1]$ since
$\frac{\mathrm{d}^{2}}{\mathrm{d}\alpha^{2}}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-e^{\gamma}\lambda_{\alpha}\left(\ln\frac{\lambda_{1}}{\lambda_{2}}\right)^{2}<0,$
hence the maximiser $\alpha^{*}$ in Definition 2.4 is unique. Differentiating (4.6) yields the critical point condition
$0=\frac{\mathrm{d}}{\mathrm{d}\alpha}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=\lambda_{1}-\lambda_{2}-e^{\gamma}\lambda_{\alpha}\ln\!\Big(\frac{\lambda_{1}}{\lambda_{2}}\Big).$
Equivalently, the (unconstrained) critical point $\alpha=\widetilde{\alpha}$ satisfies
$\lambda_{\widetilde{\alpha}}=e^{-\gamma}L(\lambda_{1},\lambda_{2}),\qquad L(\lambda_{1},\lambda_{2}):=\frac{\lambda_{1}-\lambda_{2}}{\ln\lambda_{1}-\ln\lambda_{2}}.$
In contrast to the unweighted case ($\gamma=0$), the context tilt $\gamma$ may push the optimal Chernoff parameter to the boundary $\alpha^{*}\in\{0,1\}$.
Thus, the unconstrained maximiser is
$\widetilde{\alpha}=\frac{\ln L(\lambda_{1},\lambda_{2})-\gamma-\ln\lambda_{2}}{\ln\lambda_{1}-\ln\lambda_{2}},$
[…]
$\Pi_{[0,1]}(a):=\min\{1,\max\{0,a\}\}.$
Finally,
$D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=D^{\mathrm{w}}_{B,\alpha^{*}}(\mathbb{P},\mathbb{Q}).$
If $\lambda_{1}=\lambda_{2}$, then $\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ does not depend on $\alpha$ and $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})$ for any $\alpha\in[0,1]$.
Derivation of (4.5). For $k\in\mathbb{N}_{0}$,
$p(k)^{\alpha}q(k)^{1-\alpha}=\exp\!\{-\alpha\lambda_{1}-(1-\alpha)\lambda_{2}\}\,\frac{\lambda_{\alpha}^{k}}{k!},$
so multiplying by $e^{\gamma k}$ and summing over $k$ gives
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ $=\exp\{-\alpha\lambda_{1}-(1-\alpha)\lambda_{2}\}\sum_{k\geq 0}(e^{\gamma}\lambda_{\alpha})^{k}/k!$
$=\exp\{-\alpha\lambda_{1}-(1-\alpha)\lambda_{2}+e^{\gamma}\lambda_{\alpha}\}.$
$\Box$
4.3 Exponential models
Example 4.4 (Exponential model with exponential weight).
Let ${\cal X}=\mathbb{R}_{+}=[0,\infty)$ with Lebesgue measure. Fix two hypotheses $\mathbb{P}=\mathrm{Exp}(\lambda_{1})$ and $\mathbb{Q}=\mathrm{Exp}(\lambda_{2})$ with rates $\lambda_{1},\lambda_{2}>0$, and write
$p(x)=\lambda_{1}e^{-\lambda_{1}x}{\bf 1}\{x\geq 0\},\qquad q(x)=\lambda_{2}e^{-\lambda_{2}x}{\bf 1}\{x\geq 0\}.$
Consider the exponential weight $\varphi_{\gamma}(x)=e^{\gamma x}$ with
$\gamma<\min\{\lambda_{1},\lambda_{2}\},$
so that $\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})\in(0,\infty)$ for all $\alpha\in[0,1]$. For $\alpha\in[0,1]$ set
$\lambda_{\alpha}:=\alpha\lambda_{1}+(1-\alpha)\lambda_{2}.$
(a) Weighted Bhattacharyya coefficient and Chernoff arc. A direct computation gives
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\int_{0}^{\infty}e^{\gamma x}\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}x=\frac{\lambda_{1}^{\alpha}\lambda_{2}^{\,1-\alpha}}{\lambda_{\alpha}-\gamma}.$
Hence
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\ln(\lambda_{\alpha}-\gamma)-\alpha\ln\lambda_{1}-(1-\alpha)\ln\lambda_{2}.$
Moreover, the Chernoff-tilted density $(pq)_{\alpha}$ from (3.5) is again exponential:
$(pq)_{\alpha}(x)=\frac{e^{\gamma x}p(x)^{\alpha}q(x)^{1-\alpha}}{\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})}=(\lambda_{\alpha}-\gamma)e^{-(\lambda_{\alpha}-\gamma)x}{\bf 1}\{x\geq 0\}=\mathrm{Exp}(\lambda_{\alpha}-\gamma).$
(b) Optimal Chernoff parameter. If $\lambda_{1}\neq\lambda_{2}$, then $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})$ is strictly concave on $[0,1]$, hence the maximiser $\alpha^{*}$ in Definition 2.4 is unique. Differentiating yields the critical point condition
$\frac{\lambda_{1}-\lambda_{2}}{\lambda_{\alpha}-\gamma}=\ln\!\Big(\frac{\lambda_{1}}{\lambda_{2}}\Big).$
Equivalently, the (unconstrained) critical point $\alpha=\widetilde{\alpha}$ satisfies
$\lambda_{\widetilde{\alpha}}-\gamma=L(\lambda_{1},\lambda_{2}),\qquad L(\lambda_{1},\lambda_{2}):=\frac{\lambda_{1}-\lambda_{2}}{\ln\lambda_{1}-\ln\lambda_{2}},$
so that
$\widetilde{\alpha}=\frac{\gamma+L(\lambda_{1},\lambda_{2})-\lambda_{2}}{\lambda_{1}-\lambda_{2}}.$
[…]
$D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=D^{\mathrm{w}}_{B,\alpha^{*}}(\mathbb{P},\mathbb{Q}).$
If $\lambda_{1}=\lambda_{2}=\lambda$, then $\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\lambda/(\lambda-\gamma)$ does not depend on $\alpha$, so any $\alpha\in[0,1]$ is optimal and $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\ln(\lambda-\gamma)-\ln\lambda$. Setting $\gamma=0$ (i.e. $\varphi\equiv 1$) recovers the classical unweighted expressions.
Additional example (baseline, non-exponential family).
Appendix 7 contains a closed-form illustration for the Cauchy location–scale family. Since the Cauchy family is not an exponential family, this example complements the main text by showing that, even in the unweighted baseline case $\varphi\equiv 1$, the Bhattacharyya coefficient (in particular $\rho_{1/2}$) and the Chernoff information may involve special functions (complete elliptic integrals). For nontrivial weights $\varphi$, the symmetry $\rho_{\alpha}=\rho_{1-\alpha}$ (hence $\alpha^{*}=1/2$) typically fails and a comparable closed form is not available, so we keep the baseline Cauchy computation in the appendix.
4.4 Extension to $M$-ary hypothesis testing
We can record a direct $M$-ary analogue of Theorem 3.1 for a finite family of simple hypotheses via the optimal $M$-ary pointwise loss between pairwise minima (Lemma 4.6), and each pairwise term has the exponent given by the weighted Chernoff information. Thus the $M$-ary exponent is governed by the closest pair in terms of $D_{C}^{\mathrm{w}}$.
Fix an integer $M\geq 2$ and let $\mathbb{P}_{1},\ldots,\mathbb{P}_{M}$ be probability measures on $\cal X$} dominated by $\mu$, with strictly positive densities $p_{1},\ldots,p_{M}$. Let $X_{1}^{n}=(X_{1},\ldots,X_{n})$ be i.i.d. under each hypothesis $H_{i}:\ X_{1}^{n}\sim\mathbb{P}_{i}^{\otimes n}$. Assume that the weight function factorises as in Assumption 1.1.
Assume moreover that for every $1\leq i<j\leq M$ and every $\alpha\in[0,1]$,
$\rho^{\mathrm{w}}_{\alpha}(p_{i},p_{j})=\int_{\cal X}\varphi(x)\,p_{i}(x)^{\alpha}p_{j}(x)^{1-\alpha}\,\mathrm{d}\mu(x)\in(0,\infty),$
so that all pairwise weighted Chernoff informations are well-defined.
A (deterministic) $M$-ary decision rule is a measurable map $\delta_{n}:{\cal X}^{n}\to\{1,\ldots,M\}$. Define the context-sensitive loss under $H_{i}$ by
$L_{i,n}(\delta_{n}):=\mathbb{E}_{\mathbb{P}_{i}^{\otimes n}}\!\left[\varphi(X_{1}^{n})\,{\bf 1}\{\delta_{n}(X_{1}^{n})\neq i\}\right],$
and the total loss
$L_{n,M}(\delta_{n}):=\sum_{i=1}^{M}L_{i,n}(\delta_{n}),\qquad L_{n,M}^{*}:=\inf_{\delta_{n}}L_{n,M}(\delta_{n}).$
Proposition 4.5 (Pointwise form of the optimal $M$-ary total loss).
For each $n\geq 1$,
$L_{n,M}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\left(\sum_{i=1}^{M}p_{i}(x_{1}^{n})-\max_{1\leq j\leq M}p_{j}(x_{1}^{n})\right)\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (4.7)
where $p_{i}(x_{1}^{n})=\prod_{k=1}^{n}p_{i}(x_{k})$. Moreover, an optimal rule is given by the maximum-likelihood classifier
$\delta_{n}^{*}(x_{1}^{n})\in\arg\max_{1\leq j\leq M}p_{j}(x_{1}^{n})$
(with any measurable tie-breaking).
Proof.
Fix $\delta_{n}$. Using $\sum_{i=1}^{M}{\bf 1}\{\delta_{n}\neq i\}p_{i}=\sum_{i=1}^{M}p_{i}-p_{\delta_{n}}$ pointwise, we obtain
$L_{n,M}(\delta_{n})=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\left(\sum_{i=1}^{M}p_{i}(x_{1}^{n})-p_{\delta_{n}(x_{1}^{n})}(x_{1}^{n})\right)\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$
Minimisation over $\delta_{n}$ is therefore pointwise in $x_{1}^{n}$ and is achieved by selecting an index maximising $p_{j}(x_{1}^{n})$, yielding (4.7). ∎
Lemma 4.6 (Pairwise minima).
For any nonnegative numbers $a_{1},\ldots,a_{M}$,
$\max_{1\leq i<j\leq M}\min(a_{i},a_{j})\ \leq\ \sum_{k=1}^{M}a_{k}-\max_{1\leq k\leq M}a_{k}\ \leq\ \sum_{1\leq i<j\leq M}\min(a_{i},a_{j}).$ (4.8)
Consequently, defining for $i<j$
$I_{n}^{i,j}:=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\min\{p_{i}(x_{1}^{n}),p_{j}(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$
we have the sandwich inequality
$\max_{i<j}I_{n}^{i,j}\ \leq\ L_{n,M}^{*}\ \leq\ \sum_{i<j}I_{n}^{i,j}.$ (4.9)
Proof.
Let $a_{(1)}\geq a_{(2)}\geq\cdots\geq a_{(M)}$ be the decreasing rearrangement of $(a_{1},\ldots,a_{M})$. Then $\sum_{k}a_{k}-\max_{k}a_{k}=\sum_{r=2}^{M}a_{(r)}\geq a_{(2)}$. Moreover, $\max\limits_{i<j}\min(a_{i},a_{j})=a_{(2)}$, proving the left inequality in (4.8).
For the right inequality, let $k^{*}\in\arg\max_{k}a_{k}$. Then
$\sum_{1\leq i<j\leq M}\min(a_{i},a_{j})\ \geq\ \sum_{k\neq k^{*}}\min(a_{k},a_{k^{*}})=\sum_{k\neq k^{*}}a_{k}=\sum_{k=1}^{M}a_{k}-\max_{k}a_{k}.$
Applying (4.8) pointwise to $a_{i}=p_{i}(x_{1}^{n})$, multiplying by $\varphi(x_{1}^{n})$ and integrating yields (4.9). ∎
Theorem 4.7 ($M$-ary exponent equals the minimum pairwise weighted Chernoff information).
For $1\leq i<j\leq M$, let $D_{C}^{\mathrm{w}}(\mathbb{P}_{i},\mathbb{P}_{j})$ be the weighted Chernoff information as in Definition 2.4. Set
$C_{M}^{\mathrm{w}}:=\min_{1\leq i<j\leq M}D_{C}^{\mathrm{w}}(\mathbb{P}_{i},\mathbb{P}_{j}).$
Then the optimal $M$-ary total loss satisfies
$L_{n,M}^{*}=\exp\{-nC_{M}^{\mathrm{w}}+o(n)\},\qquad n\to\infty,$ (4.10)
or equivalently,
$\lim_{n\to\infty}-\frac{1}{n}\ln L_{n,M}^{*}=C_{M}^{\mathrm{w}}.$
Proof.
Fix $1\leq i<j\leq M$ and consider the binary testing problem between $\mathbb{P}_{i}$ and $\mathbb{P}_{j}$ with the same factorised weight $\varphi$. The optimal binary total loss equals
$I_{n}^{i,j}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\min\{p_{i}(x_{1}^{n}),p_{j}(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$
and by Theorem 3.1 applied to the pair $(\mathbb{P}_{i},\mathbb{P}_{j})$,
$I_{n}^{i,j}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P}_{i},\mathbb{P}_{j})+o_{i,j}(n)\}.$
Since the number of pairs is finite, letting $r_{n}:=\max_{i<j}|o_{i,j}(n)|$ yields $r_{n}=o(n)$ and
$I_{n}^{i,j}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P}_{i},\mathbb{P}_{j})+O(r_{n})\}\qquad\text{uniformly over }i<j.$
Now use the sandwich inequality(4.9). Let $(i^{*},j^{*})$ attain the minimum $C_{M}^{\mathrm{w}}=D_{C}^{\mathrm{w}}(\mathbb{P}_{i^{*}},\mathbb{P}_{j^{*}})$. From the lower bound,
$L_{n,M}^{*}\geq I_{n}^{i^{*},j^{*}}=\exp\{-nC_{M}^{\mathrm{w}}+O(r_{n})\}.$
From the upper bound,
$L_{n,M}^{*}\leq\sum_{i<j}I_{n}^{i,j}\leq\binom{M}{2}\exp\{-nC_{M}^{\mathrm{w}}+O(r_{n})\}.$
Taking $\frac{1}{n}\ln(\cdot)$ and letting $n\to\infty$ yields (4.10). ∎
Remark 4.8 (Nonzero priors do not change the exponent).
Let $w_{1},\ldots,w_{M}>0$, $\sum_{i}w_{i}=1$, and consider the Bayesian weighted total loss $L_{n,M}^{(w)}(\delta_{n})=\sum\limits_{i=1}^{M}w_{i}L_{i,n}(\delta_{n})$ with optimum $L_{n,M}^{(w)*}:=\inf_{\delta_{n}}L_{n,M}^{(w)}(\delta_{n})$. Then the exponent remains $C_{M}^{\mathrm{w}}$. Indeed, writing $w_{\min}:=\min_{i}w_{i}$ and $w_{\max}:=\max_{i}w_{i}$, for any $\delta_{n}$,
$w_{\min}L_{n,M}(\delta_{n})\leq L_{n,M}^{(w)}(\delta_{n})\leq w_{\max}L_{n,M}(\delta_{n}),$
and taking infimum over $\delta_{n}$ gives $w_{\min}L_{n,M}^{*}\leq L_{n,M}^{(w)*}\leq w_{\max}L_{n,M}^{*}$. Hence $\frac{1}{n}\ln L_{n,M}^{(w)*}$ and $\frac{1}{n}\ln L_{n,M}^{*}$ have the same limit $C_{M}^{\mathrm{w}}$.
5 Conclusion
We studied context-sensitive (weighted) simple hypothesis testing under a multiplicative weight function $\varphi$ and showed that the optimal total loss (sum of type-I and type-II context-sensitive losses) admits an exact single-letter logarithmic asymptotic with exponent given by the weighted Chernoff information $D_{C}^{\mathrm{w}}$. A key technical step is an exponential-family embedding of the weighted geometric mixtures $\varphi\,p^{\alpha}q^{1-\alpha}$, which identifies the optimal Chernoff parameter $\alpha^{*}$ through the log-normaliser and yields information-geometric identities (including weighted Bregman representations). We also established finite-$n$ concentration bounds for the tilted weighted log-likelihood via martingale inequalities and derived explicit closed-form formulas for $D^{\mathrm{w}}_{B,\alpha}$ and $D_{C}^{\mathrm{w}}$ in several standard parametric models (Gaussian, Poisson, exponential), recovering the classical unweighted expressions when $\varphi\equiv 1$; an extension to finitely many hypotheses follows by reduction to pairwise weighted Chernoff exponents.
6 Acknowledment
The work by MK was done in the framework of the Basic Research Program of HSE.
7 Appendix: Cauchy location–scale family
We consider the univariate Cauchy location–scale family
$p_{l,s}(x)=\frac{s}{\pi\bigl(s^{2}+(x-l)^{2}\bigr)},\qquad x\in\mathbb{R},\quad l\in\mathbb{R},\ s>0.$
In this appendix we set $\varphi\equiv 1$, so the weighted quantities from Section 2 coincide with their classical counterparts (e.g. $\rho^{\mathrm{w}}_{\alpha}=\rho_{\alpha}$, $D^{\mathrm{w}}_{B,\alpha}=D_{B,\alpha}$, $D_{C}^{\mathrm{w}}=D_{C}$). This example provides a closed-form illustration outside the exponential-family setting.
Although the main body of the paper develops weighted Chernoff/Bhattacharyya quantities, we include the Cauchy location–scale family as an unweighted baseline closed-form benchmark outside the exponential-family setting (note the appearance of complete elliptic integrals even in the classical case). This appendix is not used in the proofs of the weighted results; it serves as a sanity check and illustrates the analytic complexity of $\rho_{\alpha}$ beyond exponential families. Nontrivial weights must satisfy the finiteness conditions from Section 2; for heavy-tailed Cauchy laws, common exponential tilts $\varphi(x)=e^{\gamma x}$ violate these conditions for any $\gamma\neq 0$, and in general a nonconstant weight also breaks the symmetry leading to $\alpha^{*}=1/2$.
KL divergence (closed form).
Proposition 7.1 (KL divergence between two Cauchy laws).
Let $\mathbb{P}=\mathrm{Cauchy}(l_{1},s_{1})$ and $\mathbb{Q}=\mathrm{Cauchy}(l_{2},s_{2})$ with $s_{1},s_{2}>0$. Then
$D_{\rm KL}(\mathbb{P}||\mathbb{Q})=\int_{\mathbb{R}}p_{l_{1},s_{1}}(x)\,\ln\frac{p_{l_{1},s_{1}}(x)}{p_{l_{2},s_{2}}(x)}\,\mathrm{d}x=\ln\frac{(s_{1}+s_{2})^{2}+(l_{1}-l_{2})^{2}}{4s_{1}s_{2}}.$ (7.1)
Chernoff parameter and Bhattacharyya coefficient.
Recall the (unweighted) $\alpha$-Bhattacharyya coefficient and distance
$\rho_{\alpha}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}p_{l_{1},s_{1}}(x)^{\alpha}\,p_{l_{2},s_{2}}(x)^{1-\alpha}\,\mathrm{d}x,\penalty 10000\ D_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho_{\alpha}(\mathbb{P},\mathbb{Q}),\penalty 10000\ \alpha\in[0,1].$
Proposition 7.2 (Chernoff parameter for Cauchy).
For the Cauchy location–scale family (with $\varphi\equiv 1$), one has the symmetry
$\rho_{\alpha}(\mathbb{P},\mathbb{Q})=\rho_{1-\alpha}(\mathbb{P},\mathbb{Q}),\qquad\alpha\in[0,1],$ (7.2)
equivalently $D_{B,\alpha}(\mathbb{P},\mathbb{Q})=D_{B,1-\alpha}(\mathbb{P},\mathbb{Q})$. Consequently, if $\mathbb{P}\neq\mathbb{Q}$, then the Chernoff maximiser is unique and equals
$\alpha^{*}=\frac{1}{2},\qquad\text{and hence}\qquad D_{C}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}D_{B,\alpha}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q}).$
If $\mathbb{P}=\mathbb{Q}$, then $D_{B,\alpha}(\mathbb{P},\mathbb{Q})\equiv 0$ and every $\alpha\in[0,1]$ is a maximiser.
Proof.
Set $F(\alpha):=\ln\rho_{\alpha}(\mathbb{P},\mathbb{Q})$. By Hölder’s inequality, $F$ is convex on $[0,1]$, and it is strictly convex unless $p_{l_{1},s_{1}}/p_{l_{2},s_{2}}$ is $\mu$-a.e. constant (equivalently, unless $\mathbb{P}=\mathbb{Q}$). Hence $D_{B,\alpha}(\mathbb{P},\mathbb{Q})=-F(\alpha)$ is concave (strictly concave when $\mathbb{P}\neq\mathbb{Q}$). For Cauchy laws, the symmetry (7.2) holds; see, e.g., [8] for an invariance-based proof. Therefore $D_{B,\alpha}$ is symmetric about $1/2$, and concavity implies that it attains its maximum at $\alpha=1/2$. Strict concavity yields uniqueness when $\mathbb{P}\neq\mathbb{Q}$. ∎
Closed form for $\rho_{1/2}$ and $D_{C}$.
Lemma 7.3 (A standard elliptic-integral identity).
Let $a,b>0$ and $d\in\mathbb{R}$. Define the complete elliptic integral of the first kind
$K(m):=\int_{0}^{\pi/2}\frac{\mathrm{d}u}{\sqrt{1-m\sin^{2}u}},\qquad m\in[0,1).$
Then
$\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{\sqrt{(x^{2}+a^{2})\bigl((x-d)^{2}+b^{2}\bigr)}}=\frac{4}{\sqrt{(a+b)^{2}+d^{2}}}\,K\!\left(\frac{(a-b)^{2}+d^{2}}{(a+b)^{2}+d^{2}}\right).$ (7.3)
Proposition 7.4 (Bhattacharyya coefficient for Cauchy, closed form).
Let $\mathbb{P}=\mathrm{Cauchy}(l_{1},s_{1})$ and $\mathbb{Q}=\mathrm{Cauchy}(l_{2},s_{2})$ with $s_{1},s_{2}>0$, and set $\delta:=l_{1}-l_{2}$. Then
$\rho_{1/2}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}\sqrt{p_{l_{1},s_{1}}(x)\,p_{l_{2},s_{2}}(x)}\,\mathrm{d}x=\frac{4\sqrt{s_{1}s_{2}}}{\pi\sqrt{(s_{1}+s_{2})^{2}+\delta^{2}}}\,K\!\left(\frac{(s_{1}-s_{2})^{2}+\delta^{2}}{(s_{1}+s_{2})^{2}+\delta^{2}}\right).$ (7.4)
Consequently,
$D_{C}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q})=-\ln\rho_{1/2}(\mathbb{P},\mathbb{Q}),$ (7.5)
with $\rho_{1/2}(\mathbb{P},\mathbb{Q})$ given by (7.4).
Proof.
By definition,
$\sqrt{p_{l_{1},s_{1}}(x)\,p_{l_{2},s_{2}}(x)}=\frac{\sqrt{s_{1}s_{2}}}{\pi\sqrt{\bigl((x-l_{1})^{2}+s_{1}^{2}\bigr)\bigl((x-l_{2})^{2}+s_{2}^{2}\bigr)}}.$
Shift $x\mapsto x+l_{2}$ to obtain
$\rho_{1/2}(\mathbb{P},\mathbb{Q})=\frac{\sqrt{s_{1}s_{2}}}{\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{\sqrt{(x^{2}+s_{2}^{2})\bigl((x-\delta)^{2}+s_{1}^{2}\bigr)}}.$
Apply Lemma 7.3 with $a=s_{2}$, $b=s_{1}$, $d=\delta$ to get (7.4). Finally, Proposition 7.2 yields $D_{C}(\mathbb{P},\mathbb{Q})=D_{B,1/2}(\mathbb{P},\mathbb{Q})$, hence (7.5). ∎
Remark 7.5 (Useful special cases).
If $l_{1}=l_{2}$ (common location), then $\delta=0$ and (7.4) reduces to
$\rho_{1/2}(\mathbb{P},\mathbb{Q})=\frac{4\sqrt{s_{1}s_{2}}}{\pi(s_{1}+s_{2})}\,K\!\left(\left(\frac{s_{1}-s_{2}}{s_{1}+s_{2}}\right)^{2}\right).$
If $s_{1}=s_{2}=s$ (common scale), then
$\rho_{1/2}(\mathbb{P},\mathbb{Q})=\frac{4s}{\pi\sqrt{4s^{2}+\delta^{2}}}\,K\!\left(\frac{\delta^{2}}{4s^{2}+\delta^{2}}\right).$
Remark 7.6 (Weighted case).
For a general weight $\varphi$, one has $\rho^{\mathrm{w}}_{1/2}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}\varphi(x)\sqrt{p_{l_{1},s_{1}}(x)p_{l_{2},s_{2}}(x)}\,\mathrm{d}x$. The symmetry (7.2) (and hence $\alpha^{*}=1/2$) typically fails unless $\varphi$ is compatible with the invariance argument used in [8]. We therefore use the Cauchy model mainly as a closed-form baseline example at $\varphi\equiv 1$.